OpenStudy (anonymous):

8 questions missed on Algebra 2. Help please!

OpenStudy (anonymous):

Simplify the sum. \[\frac{d^2+d-30 }{ d^2+3d-40 }\] +\[\frac{ d^2+14d+48 }{d^2-2d-48 }\]

OpenStudy (anonymous):

Simplify the complex fraction. \[ \frac{ \frac{ y-1 }{ y^2+y-6 } }{\frac{ y-6 }{ y+3 } }\]

OpenStudy (anonymous):

\[\frac{ d^2+d-30 }{ d^2+3d -40}=\frac{ (d-5)(d+6) }{(d-5)(d+8) }\] \[\frac{ d^2+14d+48 }{ d^2-2d-48 }=\frac{ (d+8) (d+6)}{(d-8)(d+6) }\]